# Out of the Memory Barrier: A Highly Memory Efficient Training System for LLMs with Million-Token Contexts

### v2

@@ 2 Related Work
Given these limitations, state-of-the-art models still rely on fine-tuning to expand their context windows (Grattafiori et al., 2024; Yang et al., 2025; Liu et al., 2025a). While recent advances in parameter-efficient fine-tuning (PEFT) have successfully minimized the memory footprint of model weights and optimizer states (Zhao et al., 2025; Zhang et al., 2025; Zhong et al., 2025; Ke et al., 2025), they do not alleviate the activation memory bottleneck that scales linearly with sequence length. Consequently, the prevailing trend has shifted from algorithmic efficiency (Chen et al., 2024) toward overcoming memory barriers with massive computational resources (Liu et al., 2024; Li et al., 2024; Jacobs et al., 2024; Shoeybi et al., 2020). For instance, recent efforts have required a 256-GPU cluster to extend a model’s context window to four million tokens (Xu et al., 2025).
@@ 4.2 Sparse Attention
The quadratic complexity of attention is a primary bottleneck in long-context training. OOMB addresses this by integrating page-level sparse attention, which reduces computational cost and is naturally supported by our paged KV cache architecture (Chen et al., 2024; Li et al., 2025; Ouyang et al., 2024). We implement two variants of sparse attention tailored to different model architectures.
Approximating Dense Attention for Qwen2.5. For models with dense attention like Qwen2.5 (Yang et al., 2025), we employ a Top-K page retrieval method to approximate the full attention pattern (Tang et al., 2024; Lu et al., 2025; Ouyang et al., 2025). The selected page indices from the forward pass are cached for reuse during backpropagation. The retrieval process begins by computing a single representative vector $K_{\text{avg}}\in\mathbb{R}^{n\times D}$ for each of the $n$ key pages by averaging. Given queries $Q\in\mathbb{R}^{(m\times P)\times D}$ from the current chunk, where $m$ is the number of query pages in the chunk and $P$ is the page size (tokens per page), we compute similarity scores between each query token and all page-representative vectors. A voting mechanism then aggregates these scores for each of the $m$ query pages to identify the most relevant key pages:

### v3

@@ 2 Related Work
Given these limitations, state-of-the-art models still rely on fine-tuning to expand their context windows (Grattafiori et al., 2024; Yang et al., 2025; Liu et al., 2025a). While recent advances in parameter-efficient fine-tuning (PEFT) have successfully minimized the memory footprint of model weights and optimizer states (Zhao et al., 2025; Zhang et al., 2025; Zhong et al., 2025), they do not alleviate the activation memory bottleneck that scales linearly with sequence length. Consequently, the prevailing trend has shifted from algorithmic efficiency (Chen et al., 2024) toward overcoming memory barriers with massive computational resources (Liu et al., 2024; Li et al., 2024; Jacobs et al., 2024; Shoeybi et al., 2020). For instance, recent efforts have required a 256-GPU cluster to extend a model’s context window to four million tokens (Xu et al., 2025).
@@ 4.2 Sparse Attention
The quadratic complexity of attention is a primary bottleneck in long-context training. OOMB addresses this by integrating page-level sparse attention, which reduces computational cost and is naturally supported by our paged KV cache architecture (Chen et al., 2024; Li et al., 2025). We implement two variants of sparse attention tailored to different model architectures.
Approximating Dense Attention for Qwen2.5. For models with dense attention like Qwen2.5 (Yang et al., 2025), we employ a Top-K page retrieval method to approximate the full attention pattern (Tang et al., 2024; Lu et al., 2025). The selected page indices from the forward pass are cached for reuse during backpropagation. The retrieval process begins by computing a single representative vector $K_{\text{avg}}\in\mathbb{R}^{n\times D}$ for each of the $n$ key pages by averaging. Given queries $Q\in\mathbb{R}^{(m\times P)\times D}$ from the current chunk, where $m$ is the number of query pages in the chunk and $P$ is the page size (tokens per page), we compute similarity scores between each query token and all page-representative vectors. A voting mechanism then aggregates these scores for each of the $m$ query pages to identify the most relevant key pages: